OpenStudy (honeymonster):

please help..find the value of the unknown in each formula.. posted in replies

OpenStudy (honeymonster):

\[\frac{ 3 }{ m } - \frac{ n }{ 4} = \frac{ 1 }{ p }\]m=2, n=3. find value of p

OpenStudy (anonymous):

p=2

OpenStudy (anonymous):

its 3/2 - 3/4

OpenStudy (anonymous):

opps.. ya @escolas

OpenStudy (anonymous):